# Metric3D v2: A Versatile Monocular Geometric Foundation Model for Zero-shot Metric Depth and Surface Normal Estimation

### v1

@@ 4.3 Ablation Study
Ablation on canonical transformation. We study the effect of our proposed canonical transformation for the input images (‘CSTM_input’) and the canonical transformation for the ground-truth labels (‘CSTM_output’). Results are reported in Tab. IX. We train the model on sampled mixed data (90K images) and test it on 6 datasets. A naive baseline (‘Ours w/o CSTM’) is to remove CSTM modules and enforce the same supervision as ours. Without CSTM, the model is unable to converge when training on mixed metric datasets and cannot achieve metric prediction ability on zero-shot datasets. This is why recent mixed-data training methods compromise learning the affine-invariant depth to avoid metric issues. In contrast, our two CSTM methods both can enable the model to achieve the metric prediction ability, and they can achieve comparable performance. Tab. I also shows comparable performance. Therefore, both adjusting the supervision and the input image appearance during training can solve the metric ambiguity issues. Furthermore, we compare with CamConvs [38], which encodes the camera model in the decoder with a 4-channel feature. ‘CamConvs’ employ the same training schedule, model, and training data as ours. This method enforces the network to implicitly understand various camera models from the image appearance and then bridges the imaging size to the real-world size. We believe that this method challenges the data diversity and network capacity, thus their performance is worse than ours.
@@
Ablation on canonical space. We study the effect of the canonical camera, i.e., the canonical focal length. We train the model on the small sampled dataset and test it on the validation set of training data and testing data. The average AbsRel error is calculated. We experiment on 3 different focal lengths, i.e., 500, 1000, 1500. Experiments show that $focal=1000$ has slightly better performance than others, see supplementary for more details.
Effectiveness of the random proposal normalization loss. To show the effectiveness of our proposed random proposal normalization loss (RPNL), we experiment on the sampled small dataset. Results are shown in Tab. X. We test on the DDAD, Lyft, DrivingStereo (DS), NuScenes (NS), KITTI, and NYUv2. The ‘baseline’ employs all losses except our RPNL. We compare it with ‘baseline + RPNL’ and ‘baseline + SSIL [27]’. We can observe that our proposed random proposal normalization loss can further improve the performance. In contrast, the scale-shift invariant loss [27], which does the normalization on the whole image, can only slightly improve the performance.
@@
TABLE XIII: Select the best joint optimizing steps for different ViT models. We find the best step varying with model size and choose the best-fitted steps according to the following experiment results. All models are trained following the settings in Tab. XI
@@ 5 Conclusion
Authors’ photographs and biographies not available at the time of submission.

### v4

@@ 4.3 Ablation Study
Ablation on canonical transformation. We examine the impact of our proposed canonical transformations for input images (CSTM_input) and ground-truth labels (CSTM_output). Results are presented in Table IX. We trained the model on a mixed dataset of 90,000 images and tested it across six datasets. A naive baseline (Ours w/o CSTM) removes the CSTM modules, enforcing the same supervision as our approach. Without CSTM, the model struggles to converge on mixed metric datasets and fails to achieve metric predictions on zero-shot datasets. This limitation is why recent mixed-data training methods often resort to learning affine-invariant depth to sidestep metric challenges. In contrast, both of our CSTM methods enable the model to attain metric prediction capabilities and achieve comparable performance. Table I confirms this comparable performance. Thus, adjusting supervision and the appearance of input images during training effectively addresses metric ambiguity issues. Additionally, we compared our approach with CamConvs [38], which incorporates the camera model in the decoder using a 4-channel feature. While CamConvs uses the same training schedule, model, and data, it relies on the network to implicitly learn various camera models from image appearance, linking image size to real-world dimensions. We believe this approach strains data diversity and network capacity, resulting in lower performance.
@@
Ablation on canonical space. We investigate the impact of the canonical camera, specifically the canonical focal length. The model is trained on a small sampled dataset and evaluated on both the training and validation sets. We calculate the average Absolute Relative (AbsRel) error for three different focal lengths: 250, 500, 1000, 1500, and 2500. Our experiments indicate that a focal length of 1000 yields slightly better performance than the others; further details can be found in the supplementary materials.
Effectiveness of the random proposal normalization loss. To demonstrate the effectiveness of our random proposal normalization loss (RPNL), we conducted experiments on a sampled small dataset, with results shown in Table 10. We tested on DDAD, Lyft, DrivingStereo (DS), NuScenes (NS), KITTI, and NYUv2. The ’baseline’ includes all losses except RPNL, which we compare to ’baseline + RPNL’ and ’baseline + SSIL [27]’. Our RPNL significantly enhances performance, while the scale-shift invariant loss [27], which normalizes the entire image, offers slight improvements.
@@
TABLE XIII: Select the best joint optimizing steps for different ViT models. We find the best step varying with model size. All models are trained following the settings in Tab. XI